# Testing identifying assumptions in Tobit Models

### v2

@@ 3.2. IV Tobit model
$P(c_{1}\leq Y,D\geq d_{1}|Z=z)$ $=$ $1-\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)$ (3.9)
$+$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
@@
$P(Y=0,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.10)
-- $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)\equiv\Phi^{2}_{d_{1},d_{0}},$
$P(Y=0,D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)\equiv\Phi^{2}_{d_{0}},$ (3.11)
$P(Y=0,D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(-\beta_{0}-\beta_{1}z)-\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.12)
@@
Remark 4 (Sharpness).
In the context of model 2.8 or 2.12, equalities 3.4-3.12 for all $c,d,z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$ are sharp in the sense that whenever they hold, it is possible to construct a vector of $(\tilde{Y},\tilde{D},\tilde{U},\tilde{V},Z)$ or equivalently $(\tilde{Y},\tilde{D},\tilde{W},\tilde{V},Z)$ that satisfies model 2.8 and 2.12, Assumptions 3 and 4, and induces the observed distribution on the data $(Y,D,Z)$. See Appendix B for the proof.
Remark 5 (Non-learnability).
@@
Then, the equalities 3.4-3.12 are sharp testable implications for the validity of the instrumental variable Tobit Model in 2.8 coupled with assumptions 3-4. They serve as the basis for the test procedure described in Section 5.
4. Identification
In this section, we provide heuristic arguments for the identification of the parameters of the respective models, expanding on Han and Vytlacil (2017) and Acerenza, Bartalotti, and Kédagni (2023).
4.1. Classic Tobit model
Equation (3.3) established that in the classic Tobit model,
$1-P(Y=0|D)=\Phi(\alpha_{0}+\alpha_{1}D).$
$\Phi^{-1}(1-P(Y=0|D))=\alpha_{0}+\alpha_{1}D.$
The coefficients have been normalized by dividing by the latent variable’s ($U$) square root of the variance, denoted by $\sigma$. Naturally, this implies that the ratio of the original parameters is identified since $P(Y=0|D=d)=\Phi(\alpha_{0}+\alpha_{1}D)=\Phi(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D)$. This allows the identification of the inverse mills ratio, which is a known function of the standard normal density, $\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D))$. Then,
$E(Y|D,Y>0)=\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D+\sigma\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D))$
Which in turn implies:
$Y=\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D+\sigma\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D))+e$
with $e$ mean independent of $D,\lambda(\cdot)$. Since $D$ is known and $\lambda(\cdot)$ was already identified then by considering the moment conditions for $Cov(.,D)$, $Cov(.,\lambda(\cdot))$ and $E(.)$ we get a three by three linear system with a unique solution for $\tilde{\alpha}_{0},\tilde{\alpha}_{1},\sigma$. In other words, the normality assumption allows identification by leveraging the relationship of the variable of interest and the outcome at the accumulation point ($Y=0$) to obtain the scaled parameters, and the same relationship when $Y>0$ to overcome the normalization. However, the reliance on the normality and linearity assumptions underlines the importance of adequately testing the model’s structure and assumptions in empirical applications.
4.2. IV Tobit model
Turning the focus to the IV-Tobit model, the first stage is readily identified from the linear structure, $\gamma_{1}=\frac{Cov(Z,D)}{Var(Z)}$, $\gamma_{0}=E(D)-\gamma_{1}E(Z)$. Now note that calculating $P(Y=0|Z)$:
$1-P(Y=0|Z)=\Phi(-\beta_{0}-\beta_{1}Z)$
Similarly to subsection 4.1, by inverting the normal CDF we obtain $-\beta_{1}=\frac{Cov(\Phi^{-1}(1-P(Y=0|Z)),Z)}{Var(Z)}$, $-\beta_{0}=E(\Phi^{-1}(1-P(Y=0|Z)))+\beta_{1}E(Z)$. The relationship between $\beta_{0},\beta_{1},\gamma_{0},\gamma_{1}$ identifies $(\alpha_{0},\alpha_{1})$. Finally, to identify $\rho$, let $s=c-\beta_{0}-\beta_{1}z$ and $t=d-\gamma_{0}-\gamma_{1}z$. Then, by using equation 3.7:
$P(Y\leq c,D\leq d|Z=z)$ $=$ $\Phi_{W,V}(s,t;\rho)$
Which by classic results of bivariate normal random variables if we differentiate with respect to $\rho$ is:
Recall the previous coefficients are normalized, a two-step procedure can be followed to identify the full original covariance matrix and the reduced form covariance matrix (Wooldridge, 2010, p. 683-84).
@@ 5. Testing procedure
To test the sharp equalities, we rewrite each of these equalities as two inequalities (Mourifié and Wan, 2017; Acerenza, Bartalotti, and Kédagni, 2023). For a concrete example, we first note that
$P(c_{1}\leq Y|D=d)=1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)\iff E[1\{c_{1}\leq Y\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]=0.$
@@
$E[1\{c_{1}\leq Y\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0,E[-1\{c_{1}\leq Y\}+1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0.$
In doing this for all the restrictions on the empirical distribution, we can implement a test relying on existing intersection bounds inferential methods such as Chernozhukov, Lee, and Rosen (2013), which is specifically suited to test conditional moment inequalities.
@@
The sharp testable equalities for every $c_{0},c_{1}\in\mathcal{Y}$ are given by equations 3.1 and 3.3. For a partition of the support of $Y$ into $K$ arbitrary chosen sets $C_{k}=(0,c_{k})$ such that $C_{k}\in C_{k+1}$, the following set of sufficient inequalities are, for some chosen values of $c_{k},c_{k+1}\in\mathcal{Y}$, related to the components of 3.1-3.3.
@@
$W_{k}=\left\{\begin{array}[]{cc}1\{Y=0\}-(1-\Phi(\alpha_{0}+\alpha_{1}D)),&\text{ if }k=0\\ 1\{c_{k}<Y\leq c_{k+1}\}-\Phi(c_{k+1}-\alpha_{0}-\alpha_{1}D)+\Phi(c_{k}-\alpha_{0}-\alpha_{1}D),&\text{ if }1\leq k<K\\ 1\{c_{K}<Y\}-(1-\Phi(c_{K}-\alpha_{0}-\alpha_{1}D)),&\text{ if }k=K\\ \end{array}\right.$
@@
where $\theta_{k}(D)$ collects all the inequalities being tested. The decision rule for the test is given by Chernozhukov, Lee, and Rosen (2013), we reject $H_{0}$ if
@@ IV Tobit model
For the instrumental variable Tobit model, the continuous support for both the outcome and treatment poses challenges to the implementation of the test. [^8] Note that a discrete treatment can also be accommodated as an intermediate step between the current derivation and the derivations from Acerenza, Bartalotti, and Kédagni (2023). The sharp testable equalities for every $c_{0},c_{1}\in\mathcal{Y}$ and $d_{0},d_{1}\in\mathcal{D}$ are given by equations 3.4-3.12.
@@
$W_{kq}=\left\{\begin{array}[]{cc}1\{Y=0\}1\{D\leq d_{0}\}-\Phi^{2}_{d_{0}},&\text{ if }k=0,q=0\\ 1\{Y=0\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{2}_{d_{q},d_{q+1}},&\text{ if }k=0,1\leq q<Q\\ 1\{Y=0\}1\{D\geq d_{Q}\}-\Phi^{2}_{d_{Q}},&\text{ if }k=0,q=Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{k+1},c_{k},d_{0}},&\text{ if }1\leq k<K,q=0\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{k+1},c_{k},d_{q+1},d_{q}},&\text{ if }1\leq k<K,1\leq q<Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{k+1},c_{k},d_{Q}},&\text{ if }1\leq k<K,q=Q\\ 1\{Y\geq c_{K}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{K},d_{0}},&\text{ if }k=K,q=0\\ 1\{Y\geq c_{K}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{K},d_{q+1},d_{q}},&\text{ if }k=K,1\leq q<Q\\ 1\{Y\geq c_{K}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{K},d_{Q}},&\text{ if }k=K,q=Q.\\ \end{array}\right.$
@@
where $\theta_{k}q(Z)$ collects all the inequalities being tested. The decision rule for the test is given by Chernozhukov, Lee, and Rosen (2013), we reject $H_{0}$ if
@@
To implement the test within the Chernozhukov, Lee, and Rosen (2013) intersection bounds inferential method, we use the CLR Stata package described in Chernozhukov et al. (2015). The parameters in the relevant model, for example, $\beta_{0},\beta_{1},\gamma_{0},\gamma_{1},\rho$, are replaced by their maximum likelihood estimators (MLE), and asymptotic validity of this “plug-in” test follows from the argument described by (Acerenza, Bartalotti, and Kédagni, 2023, Appendix B). Some additional details on the test implementation are discussed in Section 6.
@@
3. Violations due to the empirical distributions of the residuals differing from the implied distributions in certain values of the treatment (Classic Tobit) or instrument (IV Tobit), which indicate violations of the exogeneity of treatment or instrument (Kédagni and Mourifié, 2020).
@@
The proposed procedure can be extended to include exogenous covariates, $X$, within the linear index model in 2.8. Then, the testable equalities for classic and IV Tobit models can be derived with the additional conditioning on $X$. The test with covariates could be implemented by generating a partition of the covariate space, say in $J$ grids, similar to the partition of the exogenous variable used in Section 5. For every grid in the partition, one computes the test statistic 5.3 or 5.6 for the classic or the IV-tobit, respectively. Then obtain critical values that account for multiple testing via a Bonferroni correction. In particular, for every grid, set the critical value at the significance level $\frac{\alpha}{J}$, namely $\kappa_{1-\frac{\alpha}{J}}$.  [^9] A theoretically interesting approach would partition the joint support of all the exogenous variables and compute the test statistics across all parts of the grid. However, this approach would entail significant implementation challenges since the statistical package for Chernozhukov et al. (2015) allows for only one exogenous covariate. This procedure can be cumbersome when there are many covariates or when they are continuous.
Another route that is less computationally intensive follows Acerenza, Bartalotti, and Kédagni (2023). We illustrate it for the $IV$-tobit case, and a similar logic could be applied to the classic Tobit case. Thus, with covariates added to the linear index function, we have,
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\beta_{0}+\beta_{1}Z+\beta_{2}X+W\\ D&=&\gamma_{0}+\gamma_{1}Z+\gamma_{2}X+V\end{array}\right.$ (5.10)
@@
$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z,X=x)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
@@
Since the test is carried out nonparametrically, we face challenges when the dimension of $X$ is large, especially with continuous covariates. Another operational limitation is that the Stata’s clrbound package only allows for one conditioning variable at a time. For those practical reasons, we propose implementing a weaker, non-sharp, version of the testable equalities. Continuing with the example of Equality 5, we can integrate over the covariates $X$ (or alternatively, $Z$), taking advantage of the fact that $\mathbb{E}[W|Z=z,X=x]=0$ implies $\mathbb{E}[W|Z=z]=0$ and $\mathbb{E}[W|X=x]=0$ for all random variables $W$.
Implementation becomes very similar to the IV-Tobit discussed above by redefining the simplifying notation in equations (3.4)-(3.12), with the only difference being the inclusion of $X$ on the linear indexes in $\Phi_{W}(\cdot)$ and $\Phi_{W,V}(\cdot,\cdot,\cdot)$. For example,
$\Phi^{1}_{c_{1},c_{0},d_{1},d_{0}}$$\equiv$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho),$
and similarly for the other terms. Then, compute the new $W_{kq}$ in the same manner as in Section 5, and the intersection bounds framework considers the similar $2(K+1)(Q+1)$ inequalities, based on the partition of the support of $Y$ and $D$.
$\sup_{z}E[W_{kq}|Z=z]$$\leq$ 00 0
@@
Under Assumption 5, one could base the test in conditioning on a particular covariate $X_{C}$ instead of $Z$ by integrating the sharp Equality 5 over $Z$ and then implementing the intersection bounds procedure using $X_{C}$ as the sole conditioning variable. [^10] Assumption 5 constraints the relationship between the covariates and the latent error terms in such a way that one could combine the information on all $X$ and $Z$ in an index and construct non-sharp testable equalities that would hold conditional on this index. For example, having $\sup_{t}E[W_{kq}|Z+X_{1}+...+X_{C}=t]=0$, for $k=0,\dots,K;q=0,\dots,Q$.

### v3

@@ 4.2. IV Tobit model
$P(c_{1}\leq Y,D\geq d_{1}|Z=z)$ $=$ $1-\Phi_{W}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z)-\Phi_{V}(\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z)$ (4.8)
$+$ $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
@@
$P(Y=0,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$ (4.9)
- $\Phi_{W,V}(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)\equiv\Phi^{2}_{d_{1},d_{0}},$
$P(Y=0,D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)\equiv\Phi^{2}_{d_{0}},$ (4.10)
$P(Y=0,D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z)-\Phi_{W,V}(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$ (4.11)
@@
Then, the equalities (4.3)-(4.11) are sharp testable implications for the validity of the instrumental variable Tobit Model in equations (2.8) coupled with assumptions 3-4. We collect these results more formally in Theorem 2.
Theorem 2.
Suppose that the IV Tobit model (2.8) or its reduced form representation (2.12) along with assumptions 3-4 hold. Then, the parameters $\alpha_{0}$, $\alpha_{1}$, $\gamma_{0}$, $\gamma_{1}$, $\sigma^{2}_{U}$, $\rho_{UV}$ are identified, and equalities (4.3)-(4.11) must hold for all $c,d,z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$. Furthermore, these equalities are sharp, that is, whenever they hold, it is possible to construct a vector of $(\tilde{Y},\tilde{D},\tilde{U},\tilde{V},Z)$ or equivalently $(\tilde{Y},\tilde{D},\tilde{W},\tilde{V},Z)$ that satisfies model 2.8 and 2.12, Assumptions 3 and 4, and induces the observed distribution on the data $(Y,D,Z)$.
The proof for the equalities follows from the discussion above, and further details about sharpness are presented in Appendix B. They serve as the basis for the test procedure described in Section 5.
Proposition 3 (Non-learnability).
@@
Similarly to the discussion in Proposition 2, we conjecture that the approaches proposed in Theorem 2 can be adapted to models in which the researcher is willing to assume a known joint distribution for the latent errors, $(U,V)$, replacing the normal distribution, as long as the model is identified.
@@ 5. Testing procedure
To test the sharp equalities, we rewrite each of these equalities as two inequalities (mourifie_testing_2017; acerenza_testing_2023). For a concrete example, we first note that
$P(c_{1}\leq Y|D=d)=1-\Phi\left(\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)\iff E\left[1\{c_{1}\leq Y\}-1+\Phi\left(\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D\right)|D\right]=0.$
@@
$E\left[1\{c_{1}\leq Y\}-1+\Phi\left(\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D\right)|D\right]\leq 0,E\left[-1\{c_{1}\leq Y\}+1-\Phi\left(\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D\right)|D\right]\leq 0.$
In doing this for all the restrictions on the empirical distribution, we can implement a test relying on existing intersection bounds inferential methods such as chernozhukov_intersection_2013, which is specifically suited to test conditional moment inequalities.
@@
The sharp testable equalities for every $c_{0},c_{1}\in\mathcal{Y}$ are given by equations (4.1) and (4.2). For a partition of the support of $Y$ into $K$ arbitrary chosen sets $C_{k}=(0,c_{k})$ such that $C_{k}\in C_{k+1}$, the following set of sufficient inequalities are, for some chosen values of $c_{k},c_{k+1}\in\mathcal{Y}$, related to the components of equations (4.1)-(4.2).
@@
[…]
@@
where $\theta_{k}(D)$ collects all the inequalities being tested. The decision rule for the test is given by chernozhukov_intersection_2013, we reject $H_{0}$ if
@@ IV Tobit model
For the instrumental variable Tobit model, the continuous support for both the outcome and treatment poses challenges to the implementation of the test. [^9] Note that a discrete treatment can also be accommodated as an intermediate step between the current derivation and the derivations from acerenza_testing_2023. The sharp testable equalities for every $c_{0},c_{1}\in\mathcal{Y}$ and $d_{0},d_{1}\in\mathcal{D}$ are given by equations (4.3)-(4.11).
@@
[…]
@@
where $\theta_{k}q(Z)$ collects all the inequalities being tested. The decision rule for the test is given by chernozhukov_intersection_2013, we reject $H_{0}$ if
@@
To implement the test within the chernozhukov_intersection_2013 intersection bounds inferential method, we use the CLR Stata package described in chernozhukov_implementing_2015. The parameters in the relevant model, for example, $\beta_{0},\beta_{1},\gamma_{0},\gamma_{1},\rho$, are replaced by their maximum likelihood estimators (MLE), and asymptotic validity of this “plug-in” test follows from the argument described by (acerenza_testing_2023, Appendix B). Some additional details on the test implementation are discussed in Section 6.
@@
3. Violations due to the empirical distributions of the residuals differing from the implied distributions in certain values of the treatment (Classic Tobit) or instrument (IV Tobit), which indicate violations of the exogeneity of treatment or instrument (kedagni_generalized_2020).
Remark 2.
Alternative approaches to test the classic and IV Tobit could be considered. For example, an intuitive approach would be calculate the residuals $\hat{U}=Y-\hat{\alpha}_{0}-\hat{\alpha}_{1}D$ and compare its distribution to the (truncated) std. normal through one of the usual normality tests in the literature. [^10] We received this suggestion in several seminars and from anonymous referees, to whom we are thankful. One of the challenges in doing so is that the latent variable is recoverable (as estimated by the residuals) only if $Y=\alpha_{0}+\alpha_{1}D+U>0$, while for all observations for which $Y=0$ the only information available is that $\alpha_{0}+\alpha_{1}D\leq-U$. Thus, in our case, all observations that are “at the corner” would only give us information about U being below a certain truncation value, which depends on $D$ (and any other covariates included in the model, see below), which is a non-pivotal quantity and is not associated with any well-established distributional test. A second approach would follow li_nonparametric_2023 by testing for correct parametric functional forms for $E[1\{c_{0}\leq Y\leq c_{1}\}|D]$ against $\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}D)$, which is similar in spirit to our test if one incorporates all the equalities proposed in Theorem 1-2. That approach would also require partitioning the support of the variables being considered. A nonparametric test statistic could be used, following li_nonparametric_2023, the details of which are beyond the scope of this manuscript. Both approaches would present difficulties of implementation at least as important as the ones faced by proposed test based in the inferential approach by chernozhukov_implementing_2015 discussed above. While recognizing that CLR may lead to a conservative test, the relatively easy implementation of the test using available statistical packages is an attractive feature. Further research on different approaches and their properties could lead to valuable alternatives.
@@
The proposed procedure can be extended to include exogenous covariates, $X$, within the linear index model in 2.8. Then, the testable equalities for classic and IV Tobit models can be derived with the additional conditioning on $X$. The test with covariates could be implemented by generating a partition of the covariate space, say in $J$ grids, similar to the partition of the exogenous variable used in Section 5. For every grid in the partition, one computes the test statistic 5.3 or 5.6 for the classic or the IV-tobit, respectively. Then obtain critical values that account for multiple testing via a Bonferroni correction. In particular, for every grid, set the critical value at the significance level $\frac{\alpha}{J}$, namely $\kappa_{1-\frac{\alpha}{J}}$.  [^11] A theoretically interesting approach would partition the joint support of all the exogenous variables and compute the test statistics across all parts of the grid. However, this approach would entail significant implementation challenges since the statistical package for chernozhukov_implementing_2015 allows for only one exogenous covariate. This procedure can be cumbersome when there are many covariates or when they are continuous.
Another route that is less computationally intensive follows acerenza_testing_2023. We illustrate it for the $IV$-tobit case, and a similar logic could be applied to the classic Tobit case. Thus, with covariates added to the linear index function, we have,
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\tilde{\beta}_{0}+\tilde{\beta}_{1}Z+\tilde{\beta}_{2}X+\tilde{W}\\ D&=&\tilde{\gamma}_{0}+\tilde{\gamma}_{1}Z+\tilde{\gamma}_{2}X+\tilde{V}\end{array}\right$. (5.10)
@@
$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z,$ $X$ $=x)=\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z-\frac{\tilde{\beta}_{2}}{\sigma_{\tilde{W}}}x,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z-\frac{\tilde{\gamma_{2}}}{\sigma_{\tilde{V}}}x;\rho\right)$ (5.11)
- $\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z-\frac{\tilde{\beta}_{2}}{\sigma_{\tilde{W}}}x,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z-\frac{\tilde{\gamma_{2}}}{\sigma_{\tilde{V}}}x;\rho\right)$
- $\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z-\frac{\tilde{\beta}_{2}}{\sigma_{\tilde{W}}}x,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z-\frac{\tilde{\gamma_{2}}}{\sigma_{\tilde{V}}}x;\rho\right)$
$+$ $\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z-\frac{\tilde{\beta}_{2}}{\sigma_{\tilde{W}}}x,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z-\frac{\tilde{\gamma_{2}}}{\sigma_{\tilde{V}}}x;\rho\right)$
@@
Since the test is nonparametric, we face challenges when $X$ is high-dimensional, especially with continuous covariates. Another operational limitation is that the Stata’s clrbound package only allows for one conditioning variable at a time. For those practical reasons, we propose implementing a weaker, non-sharp, version of the testable equalities. Continuing with the example of Equality 5.11, we can integrate over the covariates $X$ (or alternatively, $Z$), taking advantage of the fact that $\mathbb{E}[W|Z=z,X=x]=0$ implies $\mathbb{E}[W|Z=z]=0$ and $\mathbb{E}[W|X=x]=0$ for all random variables $W$.
Implementation becomes very similar to the IV-Tobit discussed above by redefining the simplifying notation in equations (4.3)-(4.11), with the only difference being the inclusion of $X$ on the linear indexes in $\Phi_{W}(\cdot)$ and $\Phi_{W,V}(\cdot,\cdot,\cdot)$. For example, $\Phi^{1}_{c_{1},c_{0},d_{1},d_{0}}$ is given by equation (5.11) and similarly for the other terms. Then, compute the new $W_{kq}$ in the same manner as in Section 5, and the intersection bounds framework considers the similar $2(K+1)(Q+1)$ inequalities, based on the partition of the support of $Y$ and $D$.
$\sup_{z}E[W_{kq}|Z=z]$ $\leq$ $0$
@@
Under Assumption 5, one could base the test on conditioning on a particular covariate $X_{C}$ instead of $Z$ by integrating the sharp Equality (5.11) over $Z$ and then implementing the intersection bounds procedure using $X_{C}$ as the sole conditioning variable. [^12] Assumption 5 constraints the relationship between the covariates and the latent error terms in such a way that one could combine the information on all $X$ and $Z$ in an index and construct non-sharp testable equalities that would hold conditional on this index. For example, having $\sup_{t}E[W_{kq}|Z+X_{1}+...+X_{C}=t]=0$, for $k=0,\dots,K;q=0,\dots,Q$.